# A Fourth-Order Cut-cell Multigrid Method

### v2

@@ Abstract
To numerically solve a generic elliptic equation on two-dimensional domains with rectangular Cartesian grids, we propose a cut-cell geometric multigrid method that features (1) general algorithmic steps that apply to all forms of elliptic equations and all types of boundary conditions, (2) the versatility of handling both regular and irregular domains with arbitrarily complex topology and geometry, (3) the fourth-order accuracy even at the presence of ${\cal C}^{1}$ discontinuities on the domain boundary, and (4) the optimal complexity of $O(h^{-2})$. Test results demonstrate the generality, accuracy, efficiency, robustness, and excellent conditioning of the proposed method.
@@
elliptic equations, geometric multigrid methods, finite-volume methods, poised lattice generation, generating cut cells without small volumes.
@@ 1 Introduction
As far as fourth-order FV methods on irregular domains are concerned, we are only aware of the high-order EB method developed by Collela [7] and colleagues [9], in which the irregular domain is represented as $\Omega=\{\mathbf{x}:\phi(\mathbf{x})<0\}$ with $\phi$ being a smooth level set function $\mathbb{R}^{2}\rightarrow\mathbb{R}$, the domain boundary $\partial\Omega$ as $\phi^{-1}(0)$, and the unit normal vector of $\partial\Omega$ as $\mathbf{n}=\frac{\nabla\phi}{\|\nabla\phi\|}$. Assuming the existence of a flux vector $\mathbf{F}(u)$ satisfying ${\cal L}u=\nabla\cdot\mathbf{F}(u)$, they use the divergence theorem to transform the integral of ${\cal L}u$ over an irregular cut cell to those of $\mathbf{F}(u)$ over the cell faces. To further discretize (1) on a cut cell, they expand $\mathbf{F}(u)$ in its Taylor series, fit a local multivariate polynomial via weighted least squares [9, §2.3], calculate moments of monomials over the cell faces, and finally approximate the integral of $\nabla\cdot\mathbf{F}(u)$ over the cut cell with a linear combination of integrals of $\mathbf{F}(u)$ over nearby cell faces. As such, the approximation of the irregular geometry and the discretization of (1) are tightly coupled.
@@ 4.2.1 Poised lattice generation (PLG)
We calculate the integrals on regular cells by six-order recursive Gauss formulas. In contrast, the integral of a scalar function over an irregular cut cell is first converted by Green’s theorem to another integral on the boundary Jordan curve and then approximated by sixth-order Gauss formulas; see [36] for more details. Together with the explicit approximation of the boundary Jordan curve with cubic splines, this integral formulation makes our method robust in that its fourth accuracy holds even at the presence of kinks on the domain boundary.
@@ 5.3 A cut-cell V-cycle
Residual restriction to a coarse regular cell might involve both PLG fine cells and SFV fine cells; for example, the two PLG fine cells numbered #14 and #16 in Figure 4 and the two SFV cells to the right of them are covered by a PLG coarse cell, whose residual vanishes after a single pre-smoothing. Similarly, after errors on coarse PLG cells are interpolated to fine cells, those of fine PLG cells are immediately annihilated by one around of post-smoothing. In addition, each fine PLG cell is covered by some coarse PLG cell. These observations, together with the classical theory of geometric multigrid, furnish strong heuristics in supporting the convergence of the cut-cell V-cycle. They also suggest that both $\nu_{1}$ and $\nu_{2}$ be at least 1.

### v3

@@ Abstract
To numerically solve a generic elliptic equation on two-dimensional domains with rectangular Cartesian grids, we propose a cut-cell geometric multigrid method that features (1) general algorithmic steps that apply to two-dimensional constant-coefficient elliptic equations with both divergence and non-divergence forms and all types of boundary conditions, (2) the versatility of handling both regular and irregular domains with arbitrarily complex topology and geometry, (3) the fourth-order accuracy even at the presence of ${\cal C}^{1}$ discontinuities on the domain boundary, and (4) the optimal complexity of $O(h^{-2})$. Test results demonstrate the generality, accuracy, efficiency, robustness, and excellent conditioning of the proposed method.
@@
elliptic equations, finite-volume methods, geometric multigrid methods, poised lattice generation, generating cut cells without small volumes.
@@ 1 Introduction
As far as fourth-order FV methods on irregular domains are concerned, we are only aware of the high-order EB method developed by Colella [7] and colleagues [9], in which the irregular domain is represented as $\Omega=\{\mathbf{x}:\phi(\mathbf{x})<0\}$ with $\phi$ being a smooth level set function $\mathbb{R}^{2}\rightarrow\mathbb{R}$, the domain boundary $\partial\Omega$ as $\phi^{-1}(0)$, and the unit normal vector of $\partial\Omega$ as $\mathbf{n}=\frac{\nabla\phi}{\|\nabla\phi\|}$. Assuming the existence of a flux vector $\mathbf{F}(u)$ satisfying ${\cal L}u=\nabla\cdot\mathbf{F}(u)$, they use the divergence theorem to transform the integral of ${\cal L}u$ over an irregular cut cell to those of $\mathbf{F}(u)$ over the cell faces. To further discretize (1) on a cut cell, they expand $\mathbf{F}(u)$ in its Taylor series, fit a local multivariate polynomial via weighted least squares [9, §2.3], calculate moments of monomials over the cell faces, and finally approximate the integral of $\nabla\cdot\mathbf{F}(u)$ over the cut cell with a linear combination of integrals of $\mathbf{F}(u)$ over nearby cell faces. As such, the approximation of the irregular geometry and the discretization of (1) are tightly coupled.
@@ 4.2.1 Poised lattice generation (PLG)
We calculate the integrals on regular cells by six-order recursive Gauss formulas. In contrast, the integral of a scalar function over an irregular cut cell is first converted by Green’s theorem to another integral on the boundary Jordan curve and then approximated by sixth-order Gauss formulas; see [36] for more details. Together with the explicit approximation of the boundary Jordan curve with cubic splines, this integral formulation makes our method robust in that its fourth-order accuracy holds even at the presence of kinks on the domain boundary.
@@ 5.3 A cut-cell V-cycle
Residual restriction to a coarse regular cell might involve both PLG fine cells and SFV fine cells; for example, the two PLG fine cells numbered #14 and #16 in Figure 4 and the two SFV cells to the right of them are covered by a PLG coarse cell, whose residual vanishes after a single pre-smoothing. Similarly, after errors on coarse PLG cells are interpolated to fine cells, those of fine PLG cells are immediately annihilated by one round of post-smoothing. In addition, each fine PLG cell is covered by some coarse PLG cell. These observations, together with the classical theory of geometric multigrid, furnish strong heuristics in supporting the convergence of the cut-cell V-cycle. They also suggest that both $\nu_{1}$ and $\nu_{2}$ be at least 1.